OpenStudy (anonymous):

Suppose that 3 J of work is needed to stretch a spring from its natural length of 24 cm to a length of 38 cm. (a) How much work is needed to stretch the spring from 28 cm to 33 cm? (Round your answer to two decimal places) <-- got 65/98=0.66 J, incorrect (b) How far beyond its natural length will a force of 40 N keep the spring stretched? (Round your answer one decimal place.) <-- got 19.60 cm, incorrect I will definitely give a medal if somebody can help me find the correct answer. I'll try to type my steps here in a moment

OpenStudy (anonymous):

Ok so first I set 3 Joules equal to this integral: \[\huge 3= \int\limits_{0m}^{0.14m} kx dx\]

OpenStudy (anonymous):

if my memmory is ok, F of the spring is kx. \[W=\int\limits_{0}^{14} kx*dx\]

OpenStudy (anonymous):

f(x) = kx is Hooke's Law for springs, where k is the spring constant (Newton/meters here)

OpenStudy (anonymous):

Oh, forgot to pass cm to m

OpenStudy (anonymous):

first you have to find k

OpenStudy (anonymous):

I went ahead and subtracted out the differences for the following: 38 cm - 24 cm = 14 cm = 0.14 m 28 cm - 24 cm = 4 cm = 0.04 m 33 cm - 24 cm = 9 cm = 0.09 m

OpenStudy (anonymous):

so take it out of the integral: 3=k*0,14 k=3/0,14

OpenStudy (anonymous):

I got k = 10000/49 N/m

OpenStudy (anonymous):

\[\frac{k(0.14)^2}{2}\]

OpenStudy (anonymous):

the lower bound is just 0

OpenStudy (anonymous):

so integral\[\int\limits_{0}^{0,14}x dx= 1/2 x ^{2}|_{0}^{0,14}=0,14^{2}/2\]

OpenStudy (anonymous):

(3*2)/0.0196 = 306.122 ack originally I had (2*2)/(0.0196)

OpenStudy (anonymous):

0,0098

OpenStudy (anonymous):

\[W = 0,0098\int\limits_{0}^{0,05}xdx\]

OpenStudy (anonymous):

ohh sry, not from 0, from 0,04

OpenStudy (anonymous):

and to 0,09

OpenStudy (anonymous):

0.04 to 0.09, yes What's your k constant you got?

OpenStudy (anonymous):

0,0098

OpenStudy (anonymous):

I currently have ~306 N/m

OpenStudy (anonymous):

0.0098 N/m seems unreasonably small no?

OpenStudy (anonymous):

ya, but we got it right i think

OpenStudy (anonymous):

so 0,0098(0,09^2 -0,04^2/2)

OpenStudy (anonymous):

0.0098 = ? (/me looks to see where you might have gotten that value from)

OpenStudy (anonymous):

Woot I got it! 15000/49 \(\approx\) 306 N/m This means 0.99 J and 13.1 cm are the answers. I'll give you a medal regardless for being kind enough to try to help :-)

OpenStudy (anonymous):

yes, i was wrong k=306,122